OpenStudy (anonymous):

It says: the basketball coach made up a game to play where each player takes 10 shots at the basket. For every shot made, the player gains 10 points. for every basket missed, the player loses 15 points. My question is the player with the highest score sank 7 baskets and missed 3. What was the highest score?

OpenStudy (anonymous):

you have to do the math. 10*7 is 70 and 3*15 is 45 so 70-45=25

OpenStudy (anonymous):

So let's write the equation 7x-3y=? here x=10 and y = 45 so the answer will be 70-45=25

OpenStudy (anonymous):

you just made it look better than i did we still got the same answer

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

@apatel5 y= 15 not 45

OpenStudy (anonymous):

oops sorry it was a typo. But the answer is same.